the sum of two numbers is 27 and their product is 152, find the numbers
Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

If you factor 152, it's 19 times 8.
That's the answer, 8 and 19.

Answer by checkley77(12844)

x+y=27 or x=27-y
xy=152
(27-y)y=152
27y-y^2-152=0
-y^2+27y-152=0
y^2-27y+152=0
(y-19)(y-8)=0
y-19=0
y=19 ans. x=8 ans.
y-8=0
y=8 ans. x=19 ans.
Proofs:
19+8=27
27=27
8*19=152
152=152